# Birkhoff’s Theorem and Uniqueness: A Peak Beyond General Relativity

### v1

@@ III Extension beyond Quadratic Gravity
To establish the uniqueness of the Schwarzschild solution as the spherically symmetric vacuum solution of this class of modified gravity, we again consider the Einstein branch and work with the same metric ansatz as given in Eq.(1) along with Eq.(2). In the first order in $\epsilon$, we symbolically write the metric as $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon\,g_{\mu\nu}^{(1)}(t,r)$. Substituting it in the field equations and using the fact that $K_{\mu\nu}[g_{\mu\nu}^{Sch}]=0$, we get $G_{\mu\nu}[g_{\mu\nu}^{Sch}+\epsilon\,g_{\mu\nu}^{(1)}]=\mathcal{O}(\epsilon^{2})$  [^1] This is distinct from the Taylor expansion about $\epsilon=0$., which is just GR field equations with the deviated metric. However, since Birkhoff’s theorem holds in GR, we must have $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)$ till the first order in $\epsilon$, although the ADM mass will receive $\epsilon$-order correction as shown in the previous section. Then, $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon^{2}\,g_{\mu\nu}^{(2)}(t,r)$ up to second order. The above argument can be easily extended to all orders in $\epsilon$, completing the proof. Therefore, if a modified theory of gravity supports Schwarzschild metric as a vacuum solution, then Birkhoff’s theorem holds and Schwarzschild remains the unique spherically symmetric vacuum solution in the Einstein branch (but see the caveat in Sec. IV when complied with the junction condition for a star).
@@
A few important comments are in order. As a special case of the above construction, it is easy to see that GR vacuum shock wave (i.e., boosted BH solution to velocities tending to light’s speed producing a singular localized stress-energy tensor) solution is the unique Einstein branch shock in this class of modified gravity, extending Horowitz’s result beyond $\sigma$-model Horowitz and Steif (1990). Moreover, it follows that if $K_{\mu\nu}[g_{\mu\nu}^{Kerr}]=0$, then Kerr is the unique stationary and asymptotically flat vacuum BH solution in the Einstein branch of this particular class of modified gravity. Note that, unlike the spherically symmetric case, the stationarity and asymptotic flatness of Kerr does not follow from the field equations and need to be assumed.
IV Conclusion and Discussions
@@
To date, the extension of Birkhoff’s theorem beyond GR remains limited. However, the theoretical importance and observational consequences of such an extension to other theories can be hardly overstated. This work aims to bridge this gap for the first time by establishing Birkhoff’s theorem for a broad class of beyond-GR theories, including QG. Our proof is valid in the Einstein branches of these theories, which is smoothly connected to GR in the limit of vanishing coupling constants.
@@
Consequently, we demonstrate that the Schwarzschild solution remains the unique vacuum solution outside spherically symmetric objects in the Einstein branch of these theories. As in GR, the staticity and asymptotic flatness of the exterior spacetime follow as a direct consequence of spherical symmetry. This result has some profound implications. For instance, since the Schwarzschild metric has a vanishing Ricci scalar and tensor, it cannot be matched with any valid internal solution in the case of a horizonless object. Hence, the failure of junction conditions, which is now modified for beyond-GR theories Khakshournia and Mansouri (2023), suggest that the outside spacetime of a horizonless star cannot both be spherically symmetric and lie simultaneously in the Einstein branch.
@@
Possible future extensions of our work include relaxing the assumption of asymptotic flatness and studying the validity of Birkhoff’s theorem for asymptotically de Sitter or ant-de Sitter-type spacetimes. To this end, we note that our argument needs modification for incorporating cases of non-vacuum solutions and those modified theories that imply $K_{\mu\nu}[g_{\mu\nu}^{Sch}]\neq 0$. Moreover, it will be interesting to study this theorem beyond the Einstein branches of various modified gravity. We leave these novel studies for our future endeavors.
V Acknowledgment
@@
We thank Timothy Clifton for his valuable comments on an earlier version of this draft. We also appreciate the insightful discussions with Sumanta Chakraborty during the project’s initial stage. RG extends his gratitude to Sumanta Chakraborty for the invitation to the Indian Association for the Cultivation of Science (IACS) and to all members of the IACS Physics Department for their warm hospitality during his visit. The work of AC is supported by the National Postdoctoral Fellowship of SERB, ANRF, Government of India (File No.: PDF/2023/000550).

### v2

@@ 3 Extension beyond QG
To establish the uniqueness of the Schwarzschild solution as the spherically symmetric vacuum solution of this class of modified gravity, we again consider the Einstein branch and work with the same metric ansatz as given in Eq.(1) along with Eq.(4). In the first order in $\epsilon$, we symbolically write the metric as $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon\,g_{\mu\nu}^{(1)}(t,r)$. Substituting it in the field equations and using the fact that $K_{\mu\nu}[g_{\mu\nu}^{Sch}]=0$, we get $G_{\mu\nu}[g_{\mu\nu}^{Sch}+\epsilon\,g_{\mu\nu}^{(1)}]=\mathcal{O}(\epsilon^{2})$  [^4] This is distinct from the Taylor expansion about $\epsilon=0$ as explained earlier., which is just GR field equations with the deviated metric. However, since Birkhoff’s theorem holds in GR, we must have $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)$ till the first order in $\epsilon$, although the ADM mass will receive $\epsilon$-order correction as shown in the previous section. Then, $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon^{2}\,g_{\mu\nu}^{(2)}(t,r)$ up to second order. The above argument can be easily extended to all orders in $\epsilon$, completing the proof. Therefore, if a modified theory of gravity supports Schwarzschild metric as a vacuum solution, then Birkhoff’s theorem holds and Schwarzschild remains the unique spherically symmetric vacuum solution in the Einstein branch (but see the caveat in Sec. 4 when complied with the junction condition for a star).
@@
A few important comments are in order. As a special case of the above construction, it is easy to see that GR vacuum shock wave (i.e., boosted BH solution to velocities tending to light’s speed producing a singular localized stress-energy tensor) solution is the unique Einstein branch shock in this class of modified gravity, extending Horowitz’s result beyond $\sigma$-model Horowitz:1989bv. Moreover, it follows that if $K_{\mu\nu}[g_{\mu\nu}^{Kerr}]=0$, then Kerr is the unique stationary and asymptotically flat vacuum BH solution in the Einstein branch of this particular class of modified gravity. Note that, unlike the spherically symmetric case, the stationarity and asymptotic flatness of Kerr does not follow from the field equations and need to be assumed.
4 Conclusion and Discussions
@@
To date, the extension of Birkhoff’s theorem beyond GR remains limited. However, the theoretical importance and observational consequences of thoroughly investigating which modified theories comply with such a theorem cannot be overstated. This work aims to bridge this gap for the first time by establishing Birkhoff’s theorem for a broad class of beyond-GR theories, including QG. Our proof is valid in the Einstein branches of these theories, which is smoothly connected to GR in the limit of vanishing coupling constants.
@@
Consequently, we demonstrate that the Schwarzschild solution remains the unique vacuum solution outside spherically symmetric objects in the Einstein branch of these theories. As in GR, the staticity and asymptotic flatness of the exterior spacetime follow as a direct consequence of spherical symmetry. This result has some profound implications. For instance, since the Schwarzschild metric has a vanishing Ricci scalar and tensor, it cannot be matched with any valid internal solution in the case of a horizonless object. Hence, the failure of junction conditions, which is now modified for beyond-GR theories khakshournia2023art, suggest that the outside spacetime of a horizonless star cannot both be spherically symmetric and lie simultaneously in the Einstein branch.
@@
Possible future extensions of our work include relaxing the assumption of asymptotic flatness and studying the validity of Birkhoff’s theorem for asymptotically de Sitter or anti-de Sitter-type spacetimes. To this end, we note that our argument needs modification for incorporating cases of non-vacuum solutions and those modified theories that imply $K_{\mu\nu}[g_{\mu\nu}^{Sch}]\neq 0$. Moreover, it will be interesting to study this theorem beyond the Einstein branches of various modified gravity theories, particularly those involving non-topological Riemann terms, where our present method does not straightforwardly apply. A thorough investigation of such cases in the context of Birkhoff’s theorem is an important direction, but one that requires a dedicated and extensive study, which we leave for future our future endeavors.
Acknowledgements
@@
We are grateful to the anonymous referees for raising several important points during the review. We thank Timothy Clifton for his valuable comments on an earlier version of this draft. We also appreciate the insightful discussions with Sumanta Chakraborty during the project’s initial stage. RG extends his gratitude to Sumanta Chakraborty for the invitation to the Indian Association for the Cultivation of Science (IACS) and to all members of the IACS Physics Department for their warm hospitality during his visit. The work of AC is supported by the National Postdoctoral Fellowship of SERB, ANRF, Government of India (File No.: PDF/2023/000550).
Appendix A Explicit calculation up to first order in $\alpha$ and $\beta$
The $tr$-component of the field equation reads
$E_{tr}=\frac{\alpha}{r}\,\partial_{t}g^{(1)}_{(1,0)}(t,r)+\frac{\beta}{r}\,\partial_{t}g^{(1)}_{(0,1)}(t,r)=0.$ (9)
By virtue of Einstein branch, this equation must hold for all value of $\alpha$ and $\beta$. Hence, one must have,
$\partial_{t}g^{(1)}_{(1,0)}(t,r)=0=\partial_{t}g^{(1)}_{(0,1)}(t,r)~.$ (10)
Taking cognizance of Eq.(10), we can also write for the $rr$-component of field equation that
$\partial_{t}E_{rr}=\frac{1}{r}\partial_{t}\partial_{r}\left(\alpha\,f_{(1,0)}^{(1)}(t,r)+\beta\,f_{(0,1)}^{(1)}(t,r)\right)=0~.$ (11)
Again, by virtue of Einstein branch, we obtain
$f_{(1,0)}^{(1)}(t,r)=h_{(1,0)}^{(1)}(t)+j_{(1,0)}^{(1)}(r)$ (12)
$f_{(0,1)}^{(1)}(t,r)=h_{(0,1)}^{(1)}(t)+j_{(0,1)}^{(1)}(r)$ (13)
For brevity, we denote $\alpha\,h_{(1,0)}^{(1)}(t)+\beta\,h_{(0,1)}^{(1)}(t)\equiv\epsilon\,h^{(1)}(t)$. Now, by absorbing $h^{(1)}(t)$ in a rescaled time coordinate proves the Birkhoff’s theorem at the first order in $(\alpha,\beta)$.